OpenStudy (anonymous):

1. John Doe gets to go to the race track for a day. The track he is driving on is a perfect circle 750m in radius. If he is driving at 30m/s around the track, what is the centripetal acceleration experienced by the car, and what is the rotational velocity of the car? 2. Josh Beckett starts throwing his pitch; we assume that his arm moves in a perfect circular arc, and his arm is 0.7 m long. If he needs to release the ball with a tangential velocity of 35m/s what centripetal acceleration is required to achieve this?

OpenStudy (surry99):

What is the formula for centripetal acceleration?

OpenStudy (surry99):

ok...so what is the formula for centripetal acceleration? Hint: It depends on both r an v

OpenStudy (anonymous):

ma = m vt^2/r

OpenStudy (surry99):

so cancel the m's and what do you have?

OpenStudy (anonymous):

a= vt^2/r

OpenStudy (surry99):

yes, that is correct and can be used to solve these problems.

OpenStudy (anonymous):

How?

OpenStudy (surry99):

look at the first question...r= 750 m and v(tangentail) = 30 m/s therefore acceleration = v(tangential)^2/r = (30 m/s)^2/750m

OpenStudy (anonymous):

Oh. So the answer is 1.2 m/s^2?

OpenStudy (surry99):

yes....great!

OpenStudy (anonymous):

How about the rotational velocity?

OpenStudy (anonymous):

use velocity = rotational velocity * radius